# Joint Score-Threshold Optimization for Interpretable Risk Assessment Under Partial Supervision

### v1

@@ Abstract
Risk assessment tools in healthcare commonly employ point-based scoring systems that map patients to ordinal risk categories via thresholds. While electronic health record (EHR) data presents opportunities for data-driven optimization of these tools, two fundamental challenges impede standard supervised learning: (1) partial supervision arising from intervention-censored outcomes, where only extreme categories can be reliably labeled, and (2) asymmetric misclassification costs that increase with ordinal distance. We propose a mixed-integer programming (MIP) framework that jointly optimizes scoring weights and category thresholds under these constraints. Our approach handles partial supervision through per-instance feasible label sets, incorporates asymmetric distance-aware objectives, and prevents middle-category collapse via minimum threshold gaps. We further develop a CSO relaxation using softplus losses that preserves the ordinal structure while enabling efficient optimization. The framework supports governance constraints including sign restrictions, sparsity, and minimal modifications to incumbent tools, ensuring practical deployability in clinical workflows.
11footnotetext: Center for Health Systems and Policy Modeling, Department of Health Policy and Management, Johns Hopkins University, Baltimore, MD, USA. Email: fganjkh1@jhu.edu22footnotetext: Malone Center for Engineering in Healthcare, Johns Hopkins University, Baltimore, MD, USA33footnotetext: Center for Systems Science and Engineering, Department of Civil and Systems Engineering, Johns Hopkins University, Baltimore, MD, USA. Emails: espring6@jh.edu, kimia@jhu.edu44footnotetext: Department of Physical Medicine and Rehabilitation, School of Medicine, Johns Hopkins University, Baltimore, MD, USA. Email: ehoyer1@jhmi.edu55footnotetext: Johns Hopkins Hospital, Baltimore, MD, USA66footnotetext: Department of Physical Therapy, University of Nevada, Las Vegas, Las Vegas, NV, USA. Email: daniel.young@unlv.edu
@@ 1 Introduction
Across healthcare settings, risk assessment is routinely implemented via itemized, checklist-style instruments in which clinicians mark the presence, absence, or severity of predefined risk factors. Each factor contributes a predetermined point value to a total score, and pre-specified thresholds on this score map patients to ordinal risk categories (e.g., Low, Medium, High) that govern downstream clinical workflows such as monitoring frequency or preventive interventions [19, 16, 11, 3, 1]. This linear scoring combined with threshold-based categorization paradigm is widely adopted due to its transparency, auditability, and ease of operationalization [23, 25, 10]. The Johns Hopkins Fall Risk Assessment Tool (JHFRAT) exemplifies this approach, evaluating eight risk factors including age, fall history, medications, mobility, and cognition [19]. Each factor is assigned integer points (e.g., age $\geq$ 80 years: 3 points; impaired mobility: 2 points), and the total score maps to risk categories via thresholds: Low (0-5 points), Medium (6-13 points), and High ($\geq$ 14 points). These categories trigger standardized intervention bundles ranging from universal precautions to intensive monitoring and physical therapy consultations [17, 6].
The increasing availability of granular electronic health record (EHR) data presents an opportunity to optimize such tools through data-driven methods: re-estimating point weights, adjusting thresholds, and encoding deployment constraints while preserving the interpretable structure that clinicians trust [9, 20]. However, the practical realities of clinical deployment introduce two fundamental challenges that violate the assumptions of standard supervised learning approaches. The first challenge is partial supervision with selective labels. In deployment, patient outcomes and clinical interventions are inherently coupled. For fall risk assessment, we can construct high-confidence labels for extremes: patients who experience falls despite interventions (Safe-High) and patients who remain fall-free under minimal intervention (Safe-Low). However, patients receiving intermediate interventions have censored counterfactual outcomes–we cannot observe what would have occurred without intervention. This creates systematic label uncertainty for middle categories, necessitating a partial-label learning framework where each instance $i$ has a feasible set $\mathcal{S}_{i}$ rather than a single label. The second challenge is asymmetric, distance-aware misclassification costs. Operationally, under-triage (classifying a truly high-risk patient as low-risk) can result in preventable adverse events, while over-triage (classifying a low-risk patient as high-risk) merely increases resource utilization. Moreover, the severity of misclassification increases with ordinal distance: misclassifying a high-risk patient as low-risk is more consequential than misclassifying them as medium-risk. Training objectives must therefore incorporate both directional asymmetry and distance-awareness aligned with clinical priorities [7].
Definition 1 (Ordinal Risk Assessment Problem).
Given training data $\{(x_{i},\mathcal{S}_{i},w_{i})\}_{i=1}^{n}$ where $x_{i}\in\mathbb{R}^{p}$ are patient features, $\mathcal{S}_{i}\subseteq\mathcal{K}=\{1,\ldots,K\}$ are feasible label sets indicating possible risk categories, and $w_{i}>0$ are instance weights, the goal is to learn parameters $(\beta,\tau)$ where $\beta\in\mathbb{R}^{p}$ defines a linear scoring function $s(x)=\beta^{\top}x$ and $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$ are monotone thresholds, to minimize expected asymmetric ordinal loss subject to interpretability and operational constraints.
While ordinal regression methods have been extensively studied, existing approaches fail to address the unique challenges of clinical risk assessment. Classical ordinal regression methods such as the proportional odds model [14] assume parallel decision boundaries and complete supervision across all categories. Support vector approaches for ordinal regression [4] and threshold models [24] similarly require fully labeled training data. Recent extensions to deep learning for ordinal problems [15, 2] and specialized ranking losses [22] maintain these restrictive assumptions, treating all categories as equally important and all misclassification costs as symmetric. The partial label learning literature addresses scenarios where each instance is associated with a set of candidate labels [5, 27], with disambiguation methods that iteratively refine label assignments [28, 26] and loss-based methods that modify training objectives [8, 13]. However, our setting differs fundamentally in three ways: (1) feasible sets arise from systematic censoring due to clinical interventions rather than random labeling ambiguity, (2) we require preservation of ordinal structure throughout the learning process, and (3) misclassification costs are inherently asymmetric and distance-aware in healthcare applications [12]. Furthermore, while extensions like partial proportional odds models [18] relax the proportional odds assumption, they still assume complete supervision and do not address the selective labeling problem where entire categories may lack reliable labels.
@@
We introduce a constrained mixed-integer optimization framework that addresses these challenges while preserving the interpretable point-score structure essential for clinical adoption. Our approach formalizes the partial-label learning problem for ordinal risk assessment, where each instance $i$ has a feasible label set $\mathcal{S}_{i}$ encoding label uncertainty, and the optimization respects these constraints without fabricating labels for ambiguous cases. We jointly optimize $(\beta,\tau)$ to capture their interdependence under the ordinal structure. We embed misclassification costs that increase with ordinal distance and allow directional asymmetry, directly aligning the training objective with deployment priorities. When middle categories lack labeled examples, we prevent threshold collapse through minimum gap constraints derived from incumbent tools or cross-validation, maintaining clinically meaningful risk stratification. Finally, we derive a constrained score optimization (CSO) relaxation using softplus losses that preserves the asymmetric, distance-aware training signal while enabling efficient optimization for large-scale problems, extending ideas from smooth hinge losses [21] to the ordinal setting with partial supervision.
@@
The remainder of this paper is organized as follows. Section 2 presents our mixed-integer programming formulation and convex relaxation. Section 3 describes experimental evaluation on clinical risk assessment tasks. Section 4 discusses computational considerations, limitations, and deployment aspects. Section 5 concludes with implications for precision medicine.
2 Methods
2.1 Problem Formulation and Notation
@@
We consider the problem of learning a risk stratification tool that maps patient features $x\in\mathbb{R}^{p}$ to ordinal risk categories $\mathcal{K}=\{1,2,\ldots,K\}$ (e.g., Low, Medium, High for $K=3$). The tool consists of a linear scoring function $s(x)=\beta^{\top}x$ and monotone thresholds $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$ that partition the score space into $K$ categories. A patient with score $s(x)$ is assigned to category $k$ if $\tau_{k-1}<s(x)\leq\tau_{k}$, where we define $\tau_{0}=$-$\infty$ and $\tau_{K}=+\infty$ for notational convenience.
The key innovation in our formulation is handling partial supervision through feasible label sets. For each training instance $i$, rather than observing a single true label $y_{i}\in\mathcal{K}$, we have a feasible set $\mathcal{S}_{i}\subseteq\mathcal{K}$ that contains the possible true categories. This formulation naturally captures the selective labeling problem in clinical settings: we can confidently identify extreme cases ($\mathcal{S}_{i}=\{1\}$ for definitely low-risk, $\mathcal{S}_{i}=\{K\}$ for definitely high-risk) while middle-risk patients who received preventive interventions have ambiguous labels ($\mathcal{S}_{i}=\{1,2\}$ or $\mathcal{S}_{i}=\{2,3\}$ or even $\mathcal{S}_{i}=\{1,2,3\}$).
@@ 2.2 Mixed-Integer Programming Formulation
We formulate the joint learning of scoring weights $\beta\in\mathbb{R}^{p}$ and thresholds $\tau=(\tau_{1},\ldots,\tau_{K-1})$ as a mixed-integer program that handles partial supervision and asymmetric costs. The MIP approach provides exact solutions and naturally incorporates discrete constraints essential for clinical deployment.
Table 1 summarizes the decision variables and their interpretations. The key insight is to use binary indicators $y_{ik}$ to track whether instance $i$’s score exceeds each threshold $\tau_{k}$, and assignment variables $z_{ik}$ to determine the final category.
@@
$\beta$ $\mathbb{R}^{p}$ Scoring weights for features
$\tau_{k}$ $\mathbb{R}$ Threshold between categories $k$ and $k+1$
$s_{i}$ $\mathbb{R}$ Score for instance $i$: $s_{i}=\beta^{\top}x_{i}$
$y_{ik}$ $\{0,1\}$ Indicator: score $s_{i}$ exceeds threshold $\tau_{k}$
@@
$M$ $\mathbb{R}_{+}$ Big-M constant for constraint activation
$\varepsilon$ $\mathbb{R}_{+}$ Margin parameter for numerical stability
$\delta$ $\mathbb{R}_{+}$ Minimum gap between consecutive thresholds
The complete MIP formulation is:
$\min_{\beta,\tau,y,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{K}c(k,\mathcal{S}_{i})\cdot z_{ik}$ (1a)
subject to $\sum_{k=1}^{K}z_{ik}=1,\quad\forall i$ (1b)
$s_{i}-\tau_{k}\geq\varepsilon-M(1-y_{ik}),\quad\forall i,k$ (1d)
$s_{i}-\tau_{k}\leq-\varepsilon+My_{ik},\quad\forall i,k$ (1e)
$z_{i1}\leq 1-y_{i1},\quad\forall i$ (1f)
$z_{ik}\leq y_{i,k-1}-y_{ik},\quad\forall i,k\in\{2,\ldots,K-1\}$ (1g)
$z_{iK}\leq y_{i,K-1},\quad\forall i$ (1h)
$\beta\in\Omega,\quad y_{ik},z_{ik}\in\{0,1\}$ (1i)
@@
Equation (1b) ensures each instance is assigned to exactly one category. Constraint (1c) maintains threshold ordering with minimum separation $\delta$ to prevent degeneracy. Constraints (1d)-(1e) implement the big-M formulation linking scores to threshold crossings: $y_{ik}=1$ if and only if $s_{i}\geq\tau_{k}+\varepsilon$. The margin $\varepsilon>0$ ensures numerical stability and prevents instances from lying exactly on boundaries. Constraints (1f)-(1h) encode the logical relationship between threshold crossings and category assignments: instance $i$ is in category $k$ if it exceeds exactly the first k1k-1 thresholds.
The cost function $c(k,\mathcal{S}_{i})$ in objective (1a) encodes both the distance to the feasible set and directional asymmetry:
$c(k,\mathcal{S}_{i})=\min_{k^{*}\in\mathcal{S}_{i}}\ell(k,k^{*})$ (2)
This formulation assigns to each instance the cost of predicting category $k$ when the true category is the closest feasible one. For singleton feasible sets ($|\mathcal{S}_{i}|=1$), this reduces to standard supervised learning. For larger feasible sets, the optimization naturally selects the most consistent label within the feasible set.
@@
We use an asymmetric ordinal loss that captures clinical priorities:
$\ell(k,k^{*})=\begin{cases}\alpha_{\text{under}}\cdot|k-k^{*}|^{q}&\text{if }k<k^{*}\text{ (under-triage)}\\ \alpha_{\text{over}}\cdot|k-k^{*}|^{q}&\text{if }k>k^{*}\text{ (over-triage)}\\ 0&\text{if }k=k^{*}\end{cases},$ (3)
with $\alpha_{\text{under}}>\alpha_{\text{over}}$ to penalize under-triage more heavily. The exponent $q\geq 1$ controls the growth rate: $q=1$ yields linear penalty growth with ordinal distance, while $q=2$ imposes quadratic penalties for severe misclassifications. In fall risk assessment, we typically set $\alpha_{\text{under}}/\alpha_{\text{over}}\approx 3$ to reflect that missing a high-risk patient is approximately three times worse than over-treating a low-risk patient.
@@
The big-M constraints in (1d)-(1e) require careful selection of $M$ to ensure correctness while maintaining numerical stability. We set $M=2\cdot\max_{i}\|x_{i}\|\cdot B_{\beta}+B_{\tau}$ where $B_{\beta}$ is an upper bound on $\|\beta\|$ and $B_{\tau}$ bounds the threshold range. These bounds can be derived from domain knowledge (e.g., reasonable score ranges in clinical tools) or data-driven estimates. The MIP has $O(nK)$ binary variables and $O(nK)$ constraints, making it tractable for moderate-sized clinical datasets (thousands of patients) using modern solvers. For larger problems, we employ the CSO warm-start strategy described in Section 2.6.
2.3 Non-Degenerate Categories
When middle categories lack labeled examples (common when $\mathcal{S}_{i}\in\{\{1\},\{K\}\}$ for most instances), the optimization may collapse thresholds to eliminate these categories. While this might optimize the training objective, it defeats the clinical purpose of having gradated risk levels for differential intervention. To prevent this while avoiding label fabrication, we impose minimum gap constraints.
For the three-category case (Low/Medium/High), we enforce:
$\tau_{\text{High}}-\tau_{\text{Low}}\geq\Delta_{\min},$ (4)
where $\Delta_{\min}>0$ ensures the medium category maintains meaningful width. This parameter can be set through several approaches:
  1. 1.
    Incumbent-based: Use the gap from existing validated tools (e.g., JHFRAT has $\Delta_{\min}=8$ points between Low and High thresholds)
  2. 2.
    Cross-validation: On a held-out set with known middle-category examples, select $\Delta_{\min}$ that maximizes classification performance
  3. 3.
    Clinical reasoning: Set based on meaningful score differences (e.g., requiring at least 2-3 risk factors to differ between Low and High)
For $K>3$ categories, we can impose similar constraints on consecutive gaps or the total range, depending on which categories lack supervision.
2.4 Governance and Interpretability Constraints
The feasible set $\Omega$ in Model (1i) encodes interpretability and deployment requirements essential for clinical adoption, in particular:

### v2

@@ Abstract
Risk assessment tools in healthcare commonly employ point-based scoring systems that map patients to ordinal risk categories via thresholds. While electronic health record (EHR) data presents opportunities for data-driven optimization of these tools, two fundamental challenges impede standard supervised learning: (1) labels are often available only for extreme risk categories due to intervention-censored outcomes, and (2) misclassification cost is asymmetric and increases with ordinal distance. We propose a mixed-integer programming (MIP) framework that jointly optimizes scoring weights and category thresholds in the face of these challenges. Our approach prevents label-scarce category collapse via threshold constraints, and utilizes an asymmetric, distance-aware objective. The MIP framework supports governance constraints, including sign restrictions, sparsity, and minimal modifications to incumbent tools, ensuring practical deployability in clinical workflows. We further develop a continuous relaxation of the MIP problem to provide warm-start solutions for more efficient MIP optimization. We apply the proposed score optimization framework to a case study of inpatient falls risk assessment using the Johns Hopkins Fall Risk Assessment Tool.
Keywords: Clinical decision support, fall risk assessment, interpretable machine learning, mixed-integer programming, ordinal classification, risk scoring systems.
@@ 1 Introduction
Across healthcare settings, risk assessment is routinely implemented via itemized, checklist-style instruments in which clinicians mark the presence, absence, or severity of predefined risk factors. Each factor contributes a predetermined integer point value to a total score, and pre-specified integer thresholds map patient scores to ordinal risk categories (e.g., Low, Medium, High) that govern downstream clinical workflows such as monitoring frequency or preventive interventions [20, 18, 12, 5, 1]. The restriction to integer coefficients and thresholds is not merely conventional but essential for practical deployment, enabling rapid bedside assessment and facilitating clinician trust through cognitive transparency. This point-based scoring paradigm is widely adopted due to its transparency and ease of use: clinicians can compute risk scores through simple mental calculations without requiring calculators or computational tools [24, 27, 11]. However, many widely used point-based risk scoring tools were developed based on clinical reasoning rather than data-driven evidence, and warrant updating for current clinical practice and changing populations [28, 16].
The increasing availability of granular electronic health record (EHR) data presents an opportunity to optimize such tools through data-driven methods: re-estimating integer point weights, adjusting integer thresholds, and encoding deployment constraints while preserving both the interpretable structure and computational simplicity that clinicians trust [10, 21]. The practical realities of clinical deployment introduce two fundamental challenges that violate the assumptions of standard supervised learning approaches. The first challenge is a frequent lack of ground-truth labels for patients’ level of risk. In deployment, patient outcomes and clinical interventions are inherently coupled, preventing clear causal analysis between risk factors and outcomes. The second challenge is with promoting model assuredness and alignment with clinical priorities. Unconstrained supervised learning can result in unexpected outcomes that do not align with the realities of available resources or clinical objectives. Operationally, under-triage (classifying a truly high-risk patient as low-risk) can result in preventable adverse events with serious consequences, while over-triage (classifying a low-risk patient as high-risk) can result in resource inefficiency, alert fatigue, and worse patient outcomes. The relative severity of under- versus over-triage depends on the specific adverse event, available interventions, and patient population. While some existing models can balance this tradeoff via model hyperparameters, they are not equipped to safely control the resulting risk score distributions.
We introduce a constrained mixed-integer optimization framework for patient risk classification that addresses these challenges while preserving the interpretable point-score structure essential for clinical adoption. Our approach handles settings where labels exist only for extreme risk categories. In the extreme case (which applies in our clinical application), intervention-censored outcomes prevent reliable labeling of middle-risk categories entirely. We can confidently identify only the extremes: patients who experienced adverse events despite interventions (high-risk) and those who remained event-free without resource-intensive interventions (low-risk). Patients who received interventions but remained event-free have maximally uncertain labels, as it is not possible to fully determine whether interventions prevented adverse events or were unnecessarily applied. For these ambiguous cases, we do not force a label and exclude them from training. We retain these cases for validation, comparing the distributions of scores between risk categories across models.
@@
We embed misclassification costs that increase with ordinal distance and allow directional asymmetry, directly aligning the training objective with deployment priorities. When middle categories lack labeled examples, we prevent threshold collapse through minimum gap constraints derived from incumbent tools or cross-validation, maintaining clinically meaningful risk stratification. Finally, we extend ideas from smooth hinge losses [23] to the ordinal setting with our derived constrained score optimization (CSO) relaxation, which uses softplus losses to preserve the asymmetric, distance-aware training signal while enabling efficient optimization for large-scale problems. We develop a joint model optimization pipeline with MIP and CSO, and apply this methodology to a case study of improving the Johns Hopkins Fall Risk Assessment Tool (JHFRAT) for assessing inpatient fall risk.
@@
The remainder of this paper is organized as follows. Section 2 highlights related work in the area of ordinal risk assessment with regression methods and constrained optimization. Section 3 presents our mixed-integer programming formulation and convex relaxation. Section 4 describes experimental evaluation on clinical risk assessment tasks. Section 6 discusses computational considerations, limitations, and deployment aspects. Section 7 concludes with implications for precision medicine.
2 Related Work
Ordinal regression methods have been extensively studied, but existing approaches fail to address the unique challenges of clinical risk assessment. Classical ordinal regression methods such as the proportional odds model [15] assume parallel decision boundaries and complete supervision across all categories. Support vector approaches for ordinal regression [6] and threshold models [25] similarly require fully labeled training data. Recent extensions to deep learning for ordinal problems [17, 4] and specialized ranking losses [22] maintain these restrictive assumptions, treating all categories as equally important and all misclassification costs as symmetric. The partial label learning literature addresses scenarios where each instance is associated with a set of candidate labels [7, 30], with disambiguation methods that iteratively refine label assignments [31, 29] and loss-based methods that modify training objectives [9, 14]. However, our setting differs fundamentally in three ways: (1) feasible sets arise from systematic censoring due to clinical interventions rather than random labeling ambiguity, (2) misclassification costs are inherently asymmetric and distance-aware in healthcare applications, and (3) we require preservation of ordinal structure throughout the learning process [13]. Furthermore, while extensions like partial proportional odds models [19] relax the proportional odds assumption, they still assume complete supervision and do not address the selective labeling problem where entire categories may lack reliable labels.
We build upon the literature for constrained mixed integer optimization methods as an alternative to ordinal regression. Early work employing integer programming methods for classification includes that of Uney and Turkay [26] and Bertsimas and Shioda [2]. Both works solve supervised multi-class classification problems by separating different classes into polyhedral regions of the data space. Bertsimas and Shioda also solve a separate linear regression problem, which employs unsupervised classification to create different linear models across heterogeneous subpopulations. Billiet et al. [3] formulate a risk-score weighting binary classification linear programming model based on soft-margin support vector machines. The Interval Coded Scoring linear program presented in that work utilizes similar model constraints as integer programming classification, but avoids the use of integer or binary variables via the SVM-based objective. Dedieu et al. [8] address the computational complexity of scaling integer programming classification models by solving a series of smaller relaxed MIP subproblems. Ustun and Rudin [27] present the RiskSlim model, which optimizes integer risk model coefficients. They present both a nonlinear objective model and a surrogate linear program model. Among the current literature, their work is the most similar to our methodology, though their model applies only to binary risk classification. Furthermore, we use a discrete objective function to enforce class-based ordinal misclassification.
3 Proposed Models
@@
We consider the problem of learning a risk stratification tool that maps patient features $x\in\mathbb{R}^{p}$ to ordinal risk categories $\mathcal{K}=\{1,2,\ldots,K\}$ (e.g., Low, Medium, High for $K=3$). Following the clinical scoring system paradigm [27], the tool consists of integer-valued scoring weights $\beta\in\mathbb{Z}^{p}$ and integer thresholds $\tau=(\tau_{1},\ldots,\tau_{K-1})\in\mathbb{Z}^{K-1}$ with $\tau_{1}<\tau_{2}<\cdots<\tau_{K-1}$. The integer restriction enables clinicians to compute risk scores $s(x)=\beta^{\top}x$ through mental arithmetic without computational aids, a critical requirement for bedside deployment and cognitive transparency. A patient with score $s(x)$ is assigned to category $k$ if $\tau_{k-1}<s(x)\leq\tau_{k}$, where we define $\tau_{0}=$-$\infty$ and $\tau_{K}=+\infty$ for notational convenience.
Given training data $\{(x_{i},k_{i}^{*},w_{i})\}_{i\in\mathcal{I}}$ where $x_{i}\in\mathbb{R}^{p}$ are patient features, $k_{i}^{*}\in\mathcal{K}$ are known category labels, and $w_{i}>0$ are instance weights, our goal is to learn integer parameters $(\beta,\tau)\in\mathbb{Z}^{p}\times\mathbb{Z}^{K-1}$ that minimize expected asymmetric ordinal loss subject to interpretability and operational constraints. Our setting considers the possibility that some patient samples will not have associated risk labels. Without imposition of assumptions or causal analysis of the effects of interventions, these samples must be excluded from training to maintain the integrity of the training labels. The key challenge then becomes partial supervision at the category level: while each included training instance $i\in\mathcal{I}$ has a single known label $k_{i}^{*}$, certain categories $k\in\mathcal{K}$ may have no labeled training examples. Let $\mathcal{I}_{k}=\{i\in\mathcal{I}:k_{i}^{*}=k\}$ denote the set of training instances with known label $k$. In the extreme case where only the most extreme categories have labeled examples (i.e., $\mathcal{I}_{k}=\emptyset$ for intermediate categories), the optimization may collapse thresholds to eliminate unlabeled categories. Section 3.2 addresses this degeneracy through minimum threshold gap constraints.
@@ 3.1 Mixed-Integer Programming Formulation
We formulate the joint learning of integer scoring weights $\beta\in\mathbb{Z}^{p}$ and integer thresholds $\tau=(\tau_{1},\ldots,\tau_{K-1})\in\mathbb{Z}^{K-1}$ as a mixed-integer program that handles partial supervision and asymmetric costs. The MIP approach provides exact solutions and naturally incorporates both the integrality constraints and other discrete constraints essential for clinical deployment.
Table 1 summarizes the decision variables and parameters in our formulation. The key insight is to use assignment variables $z_{ik}$ to determine the category assignment for each instance, with big-M constraints enforcing that assigned categories respect the threshold boundaries.
The complete MIP formulation is given by:
$\min_{\beta,\tau,z}\quad$ $\sum_{i\in\mathcal{I}}w_{i}\sum_{k=1}^{K}\ell(k,k_{i}^{*})\cdot z_{ik}$ (1a)
subject to $\sum_{k=1}^{K}z_{ik}=1,\quad\forall i\in\mathcal{I}$ (1b)
$s_{i}=\beta^{T}x_{i}\quad\forall i\in\mathcal{I}$ (1d)
$M(z_{ik}-1)+\tau_{k-1}+\varepsilon\leq s_{i},$
$\qquad\forall i\in\mathcal{I},\,k\in\{1,\ldots,K\}$ (1e)
$s_{i}\leq M(1-z_{ik})+\tau_{k},$
$\qquad\forall i\in\mathcal{I},\,k\in\{1,\ldots,K\}$ (1f)
$\beta\in\Omega\cap\mathbb{Z}^{p}$ (1g)
$z\in\Theta\cap\{0,1\}^{K|\mathcal{I}|}$ (1h)
$\tau\in\mathbb{Z}^{K-1}$ (1i)
@@
$\beta$ $\mathbb{Z}^{p}$ Scoring weights
$\tau_{k}$ $\mathbb{Z}$ Threshold between categories $k$ and $k+1$
$s_{i}$ $\mathbb{Z}$ Score for instance $i$: $s_{i}=\beta^{\top}x_{i}$
@@
$M$ $\mathbb{R}_{+}$ Big-M constant
$\varepsilon$ $\mathbb{R}_{+}$ Margin parameter
$\delta$ $\mathbb{Z}_{+}$ Min. gap between thresholds
@@
Equation (1b) ensures each instance is assigned to exactly one category. Constraint (1c) maintains threshold ordering with minimum separation $\delta>0$ to prevent degeneracy, as detailed in Section 3.2. Constraints (1e)–(1f) implement the big-M formulation enforcing threshold bounds to assigned categories: if $z_{ik}=1$, then $\tau_{k-1}+\varepsilon\leq s_{i}\leq\tau_{k}$. The margin $\varepsilon>0$ ensures numerical stability and prevents ambiguity when scores lie exactly on a boundary. The domain constraints in (1g) and (1h) encapsulate governance and performance constraints on $\beta$ and $z$, respectively, that are described in Sections 3.3 and 3.4.
@@
We use an asymmetric ordinal loss function $\ell(k,k^{*})$ in objective (1a) that captures the differential consequences of misclassification:
$\ell(k,k^{*})=\begin{cases}\omega_{\text{under}}\cdot|k-k^{*}|^{q}&\text{if }k<k^{*}\\ \omega_{\text{over}}\cdot|k-k^{*}|^{q}&\text{if }k>k^{*}\\ 0&\text{if }k=k^{*}\end{cases},$ (2)
where $\omega_{\text{under}}>\omega_{\text{over}}$ reflects that under-triage typically carries more severe consequences than over-triage in healthcare settings. The exponent $q\geq 1$ controls the growth rate: $q=1$ yields linear penalty growth with ordinal distance, while $q=2$ imposes quadratic penalties for severe misclassifications. The specific ratio $\omega_{\text{under}}/\omega_{\text{over}}$ should be determined based on the clinical context, weighing the relative costs of missed adverse events versus unnecessary interventions.
@@
The big-M constraints in (1e)–(1f) require careful selection of $M$ to ensure correctness while maintaining numerical stability. We set $M=2\cdot\max_{i}\|x_{i}\|\cdot B_{\beta}+B_{\tau}$ where $B_{\beta}$ is an upper bound on $\|\beta\|$ and $B_{\tau}$ bounds the threshold range. These bounds can be derived from domain knowledge (e.g., reasonable score ranges in clinical tools) or data-driven estimates. The MIP has $O(nK)$ binary variables and $O(nK)$ constraints, making it tractable for moderate-sized clinical datasets (thousands of patients) using modern solvers. For larger problems, we employ the CSO warm-start strategy described in Section 3.6.
3.2 Handling Missing Category Labels
In the extreme partial supervision setting, where certain categories lack any labeled examples, the optimization faces a fundamental challenge: without training instances to anchor middle categories, the objective may collapse thresholds to eliminate these categories entirely. While this maximizes performance on labeled extremes, it defeats the clinical purpose of having gradated risk levels that guide differential interventions. Consider the common scenario where only extreme cases can be reliably labeled: $k^{*}_{i}\in\{1,K\}$ for all training instances. An optimizer focused solely on separating these extremes might set $\tau_{1}\approx\tau_{2}\approx\cdots\approx\tau_{K-1}$, effectively collapsing all middle categories. This would achieve optimal separation on training data but produce a binary classifier rather than the desired ordinal tool.
Therefore, to maintain clinically meaningful risk stratification without fabricating labels for unlabeled categories, we impose minimum gap constraints per (1c), where $\delta>0$ ensures all categories maintain meaningful width. This parameter can be set through several approaches. In the case where we are optimizing an existing tool, we can utilize the existing tool’s threshold gaps. For example, if the incumbent assessment tool has a gap of 8 points between low-risk and high-risk thresholds, we set $\delta=8$. On a held-out set with known middle-category examples, we can select $\delta$ via cross-validation to maximize classification performance. Alternatively, we can apply clinical reasoning to set meaningful score differences (e.g., requiring at least 2–3 risk factors to differ between Low and High categories). For $K>3$ categories, we can impose similar constraints on consecutive gaps or the total range, depending on which categories lack supervision.
3.3 Governance Constraints